# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Take a ticket and compress it between two of your fingers. It will spontaneously lose its stability and buckle left or right. Now try to push the buckled state inwards with your other hand. It will first resist, but then suddenly snap to the other side. Such buckling and snapping is a quintessential example of a critical bifurcation and has many uses: from carnivorous plants and beetles [1, 2] to MEMS [3], soft robots [4] and mechanical metamaterials [5, 6, 7, 8], buckling endows slender objects with sensitivity to environmental perturbations, multistable response and the capacity for large-scale shape morphing. Buckling and snapping are usually a single affair—they only happen if you push the beam. If instead one wishes to obtain persistent oscillations to e.g., perform work via cyclic shape changes, then one would need to constantly energize the beam from within.
@@
Here we energize buckling beams by breaking the usually reciprocal link between bending modes [16, 17, 18, 19, 20], and discover a new buckling scenario that enables tuneable snapping of beams in response to environmental perturbations. This scenario is characterised by two ingredients: activity and mechanical criticality. We show experimentally and theoretically that reciprocity-breaking leads to non-Hermitian flexural waves that advect curvature unidirectionally. This non-Hermitian dynamics modifies the critical point underpinning buckling, leading to a bifurcation mediated by a critical exceptional point (CEP) which is reminiscent of the exceptional physics of driven-dissipative condensed matter systems [21, 22, 23, 24]. By actuating our beams from within, we bypass the need for external forcing [25, 26, 27, 28, 29, 30, 31], and demonstrate crawling, walking and digging dynamics in a single free-standing filament.
Our beams are made by chaining together mechatronic linkages [Fig. 1(a), SM §V]. At the level of a single linkage, conventional bending elasticity is encoded in a torque-angle relation $\tau_{i}=-B\delta\theta_{i}$ at site $i$. By contrast, our linkages implement a torque-angle relation that couples neighboring angles antisymmetrically: $\tau_{i}=k^{\mathrm{o}}(\delta\theta_{i+1}-\delta\theta_{i-1})$, where $k^{\mathrm{o}}$ quantifies the microscopic non-reciprocity. This antisymmetric relation cannot be derived from a potential energy [32], and injects energy into the beam in a manner that breaks parity symmetry whilst conserving linear and angular momentum.
Figure 1: Non-reciprocal Buckling. Adding non-reciprocity to the canonical Euler buckling scenario causes beams to lose stability and transition into self-snapping. (a) Within a building block, each linkage has a torque-angle relationship $\tau_{i}=k^{\mathrm{o}}(\delta\theta_{i+1}-\delta\theta_{i-1})$, and responds antisymmetrically when perturbed from the left vs. right. Buckling a chain of non-reciprocal linkages under clamped tangent boundary conditions, we find that as non-reciprocity $k^{\mathrm{o}}$ increases, the beam polarizes but retains stability. Increasing activity still further, this stable state disappears and the beam persistently self-oscillates. Traces indicate linkage positions over time. Scale bar $7.5\mathrm{cm}$. (b) Decomposing the horizontal position of each linkage into normalized Fourier modes $a_{k}$, we find that the lowest two modes dominate the snapping process, oscillating out of phase with one another. Modes $a_{k}$ for $k>2$ are shown in gray, and remain negligible during snapping. (c) Smooth waves of odd torques run down the filament during snapping. (d) A single snap proceeds via a pulsation traveling from the tail of the filament to its head. Schematic indicates the torque dipoles exerted by each motor, alongside the force on the central linkage.
@@
We buckle an open segment of our active beam under clamped tangent boundary conditions with a compression $\mathrm{d}l$ [Fig. 1(a), SM §VI]. As we increase the activity $k^{\mathrm{o}}$, the initially symmetric buckled state polarizes, leaning to one side but remaining static. However, above a critical $k^{\mathrm{o}}$, this static shape loses stability and a persistent self-snapping sets in. Without a source of viscous damping within the hinges, many modes participate in this snapping dynamics [Supplementary Video 1]. Yet by increasing hinge dissipation, we drive the chain into a long-wavelength regime, where snapping is dominated only by the lowest Fourier modes [Fig. 1(b)]. Visualizing the non-reciprocal torque distribution, we find smooth stress waves that pulse down the chain in a direction determined by $k^{\mathrm{o}}$, snapping the filament back and forth [Fig. 1(c,d)]. Upon removing compression, snapping ceases and the beam stabilizes. We conclude that the nonlinear oscillations we observe are being driven by a combination two phenomena: polarization induced by non-reciprocity and buckling induced by confinement.
@@
Figure 2: Breaking reciprocity in slender filaments creates one-way flexural waves. (a, b) Perturbing an open chain of non-reciprocal linkages, we observe one-way advection of curvature. Panel (a) shows key frames of a perturbation applied at one free end. Panel (b) shows a perturbation applied at the chain midpoint, confirming the asymmetry of wave propagation. Colors indicate angular deviations from the flat state. (c) Non-reciprocal forces on a sinusoidal filament show a $\pi/2$ phase lag with respect to the filament geometry itself, leading to an antisymmetric flexural wave dispersion relation. Here $\lambda(q)$ is the complex dispersion from Eq. (2), with $\mathrm{Im}(\lambda)$ representing wave propagation and $\mathrm{Re}(\lambda)$ giving growth or decay. Dispersion shown for $\beta=\eta=1$.
@@
First we focus on the polarization of the beam by removing external confinement and studying the linear response of a stable filament. We apply a perturbation to its left edge [Fig. 2(a), SM §VI]. In a passive beam, waves would propagate and damp equally in both directions. Yet here we observe that the non-reciprocity $k^{\mathrm{o}}$ picks a clear direction of wave motion along the filament [Fig. 2(b), Supplementary Video 2]: The polarization of our clamped beam is being driven by one-way wave advection of curvature.

### v2

@@ Abstract
Figure 1: Non-reciprocal Buckling. Adding non-reciprocity to the canonical Euler buckling scenario causes beams to lose stability and transition into self-snapping. (a) Within a building block, each linkage has a torque-angle relationship $\tau_{i}=k^{\mathrm{o}}(\delta\theta_{i+1}-\delta\theta_{i-1})$, and responds antisymmetrically when perturbed from the left vs. right. (b) Buckling a chain of non-reciprocal linkages under clamped tangent boundary conditions, we find that as non-reciprocity $k^{\mathrm{o}}$ increases, the beam polarizes but retains stability. Increasing activity still further, this stable state disappears and the beam persistently self-oscillates. Traces indicate linkage positions over time. Scale bar 7.5cm7.5\ $\mathrm${cm}. (c) Decomposing the horizontal position of each linkage into normalized Fourier modes $a_{k}$, we find that the lowest two modes dominate the snapping process, oscillating out of phase with one another. Modes $a_{k}$ for $k>2$ are shown in gray, and remain negligible during snapping. (d) Smooth waves of odd torques run down the filament during snapping. (e) A single snap proceeds via a pulsation traveling from the tail of the filament to its head. The schematic indicates the torque dipoles exerted by each motor, alongside the force on the central linkage.
Take a paper ticket and compress it between two of your fingers. It will spontaneously lose its stability and buckle one way or the other. Now try to push the buckled state inwards with your other hand. It will first resist, but then suddenly snap to the other side. Such buckling and snapping is a quintessential example of a critical bifurcation and has many uses: from carnivorous plants and beetles [56, 19] to micro-electromechanical systems [4], soft robots [58] and mechanical metamaterials [12, 17, 47, 15], buckling endows slender objects with sensitivity to environmental perturbations, multistable response and the capacity for large-scale shape morphing. Buckling and snapping typically occur once—they only happen if you push the beam. If instead one wishes to obtain persistent oscillations to e.g., perform work via cyclic shape changes, then one would need to constantly energize the beam from within.
@@
Here we show that active beams with non-reciprocal internal mechanics buckle via a critical exceptional point (CEP) [3, 50, 21, 26]. This scenario is characterized by two ingredients: the mechanical criticality of buckling plus non-Hermitian linear response [8, 9, 20, 41, 33]. In conventional buckling, a single mode of a stressed beam becomes unstable via a critical point. By contrast, in CEP-mediated buckling two anti-symmetrically coupled modes of an active beam simultaneously destabilize. The resulting run-and-chase dynamics between these modes produces large oscillations in beam shape. We experimentally demonstrate that these stable oscillations are limit cycles, and develop a theoretical model that captures the nonlinear dynamics of the beam. These limit cycles are distinct from a transition to linear oscillations across exceptional points [39], because these cycles arise from the supercritical instability and post-bifurcation nonlinear dynamics of beam buckling. Finally, we harness these dynamics for multiple robotic functionalities—we demonstrate crawling, walking and digging in a single free-standing filament, without the need for external forcing [46, 11, 18, 10, 60, 48, 57]. Our beams are made by chaining together mechatronic linkages [Fig. 1(a), Methods §4]. At the level of a single linkage, conventional bending elasticity is encoded in a torque-angle relation $\tau_{i}=-B\delta\theta_{i}$ at site $i$. In contrast, our linkages implement a torque-angle relation that couples neighboring angles antisymmetrically: $\tau_{i}=k^{\mathrm{o}}(\delta\theta_{i+1}-\delta\theta_{i-1})$, where $k^{\mathrm{o}}$ quantifies the microscopic non-reciprocity. This antisymmetric relation cannot be derived from a potential energy [53], and injects energy into the beam in a manner that breaks parity symmetry whilst conserving linear and angular momentum. Whilst energy is injected into the beam by non-reciprocal couplings, it is dissipated via damping at the linkage hinges. We control dissipation by attaching a viscoelastic polymer skeleton to the backbone of the linkage. This skeleton provides bending rigidity $B$ but also acts as a torsional dashpot on each hinge, $\tau_{i}=-\Gamma\delta\dot{{\theta}}_{i}$, with a damping $\Gamma$—we calibrate these values by fitting to the dynamics of a damped torsional oscillator [Methods §IV, Figs. 7, 8].
@@
We buckle an open segment of our active beam under clamped tangent boundary conditions with a compression $\mathrm{d}l$ [Fig. 1(b), Methods Fig. 6, Supplementary Video 1]. As we increase the activity $k^{\mathrm{o}}$, the initially symmetric buckled state polarizes, leaning to one side but remaining static. However, above a critical $k^{\mathrm{o}}$, this static shape loses stability and a persistent self-snapping sets in. Without a source of damping within the hinges, many modes participate in this snapping dynamics. Yet by increasing dissipation via our viscoelastic skeleton [using Stratasys Agilus 30, calibration in Figs. 7,8], we drive the chain into a long-wavelength regime, where snapping is dominated only by the lowest Fourier modes [Fig. 1(c)]. Visualizing the non-reciprocal torque distribution, we find smooth stress waves that pulse down the chain in a direction determined by $k^{\mathrm{o}}$, snapping the filament back and forth [Fig. 1(d, e)]. Upon removing compression, snapping ceases and the beam stabilizes. We conclude that the nonlinear oscillations we observe are being driven by a combination of two phenomena: polarization induced by non-reciprocity and buckling induced by confinement.
@@
Figure 2: Breaking reciprocity in slender filaments creates one-way flexural waves. (a, b) Perturbing an uncompressed, loosely confined chain of non-reciprocal linkages, we observe one-way advection of curvature. Panel (a) shows key frames of a perturbation applied at one free end. Panel (b) shows a perturbation applied at the chain midpoint, confirming the asymmetry of wave propagation. Colors indicate angular deviations from the flat state. (c) Non-reciprocal forces on a sinusoidal filament show a $\pi/2$ phase lag with respect to the filament geometry itself, leading to an antisymmetric flexural wave dispersion relation. Here $\lambda(q)$ is the complex dispersion from (3), with $\mathrm{Im}(\lambda)$ representing wave propagation and $\mathrm{Re}(\lambda)$ giving growth or decay. Dispersion shown for $\beta=\eta=1$.
@@
First, we focus on the polarization of the beam by removing external compression and studying the linear response of a stable filament. We loosely confine a beam under zero compression, and apply a perturbation to its left edge [Fig. 2(a), Methods §3]. In a passive beam, waves would propagate and damp equally in both directions. Yet here we observe that the non-reciprocity $k^{\mathrm{o}}$ picks a clear direction of wave motion along the filament [Fig. 2(b), Supplementary Video 2]: The polarization of our clamped beam is being driven by one-way wave advection of curvature.